OpenStudy (anonymous):

write an equation of the line containing the given point & perpendicular to the given line. (0,6); 3x+7y=5

OpenStudy (anonymous):

okey dokey, rewrite the equation in y = mx+b form

OpenStudy (anonymous):

7y=5-3x y = -3x/7 + 5/7

OpenStudy (anonymous):

The slope is the coefficient of the x term, -3/7 The slope of the perpendicular line is the negative reciprocal. This means flip and negate the fraction. The slope of the perpendicular line is thus 7/3

OpenStudy (anonymous):

so the line will be: y = 7/3(x-0)+6 = 7x/3 + 6